# Useful entanglement can be extracted from noisy graph states

### v1

@@ 1 Introduction
The discovery of cluster states – states of qubits with grid-like entanglement structure – provided a new perspective for quantum computation, better applicable for some systems [1, 2]. The original design consisted of qubits arranged in a two-dimensional square lattice structure, where the neighboring qubits would be entangled by means of a controlled-$Z$ gate (or a specific quantum system equivalent thereof). Unlike the circuit-based understanding of quantum computation, the input state is constant, and the only available class of operations is the sequential measurement of individual qubits. It is the act of measurement that feeds the information into the system and performs all of the needed transformations [3]. Generalizations into different connectivity structures – the (hyper)graph states – soon followed, sharing the same core properties with the original cluster states [4, 5, 6, 7] and providing a framework for further research into the entanglement properties of multiqubit states [8, 9].
@@
The most basic operation in this context is no processing at all: a simple transfer or teleportation [10, 11] of one qubit state into another place. It is a basic building block of measurement-based quantum computation, on top of which more involved operations are built. In the basic scenario, this can be performed by sequential measurement of Pauli-$X$ operators on a path between the initial and end qubits (the terminal qubits in this article nomenclature) and Pauli-$Z$ on the qubits neighboring the path. Intuitively, the $Z$-basis measurement excises the entanglement contained in the path from the rest of the cluster state and the $X$-basis measurement along the path makes the terminal qubits correlated: final measurement of one of them fuses the two together, and the surviving qubit shares correlations with a distant part of the graph [3]. This process can also be understood as a Bell pair generation along the path: local operations do commute, and a measurement along the path on its own does produce Bell-like correlations across the terminal qubits, which can be used directly for information transfer. All of this works, provided that no noise was present during the state preparation and measurement – but this is rarely the case. Thus, a noise-detecting or -correcting process is needed, and research in this direction has been done [12, 13, 14, 15, 16].
@@
In general, entanglement is required for quantum computation, and preparation of a Bell pair or GHZ state is an operational primitive of many quantum algorithms. Extracting those states from larger graph states was discussed in [17, 18, 14, 19, 20].
@@
In this article we analyze experimentally motivated noise models and determine what can be done (to the initial noisy graph state structure and the measurement pattern) in order to mitigate the noise effects. This approach does not translate into full-fledged error correcting codes, but still uses the properties of the stabilizer associated with the graph state; in the ideal case, some outcomes are perfectly (anti)correlated, and any deviation indicates the presence of noise. As mentioned in the previous paragraph, teleportation is equivalent to a Bell pair preparation; in order to simplify the calculations, we analyze the entanglement quality of the prepared state after the internal qubits have been measured.
@@
The article is structured as follows. In Section 2 we briefly introduce graph states and the stabilizer formalism. In Section 3 we explain the measurement scheme which we use to generate Bell pairs between certain nodes or for teleportation. We introduce noise models in Section 4 and analyze how to deal with their effects in Section 5. In Section 6 we show how to extend the presented methods to generate GHZ states. We conclude our work in Section 7.
@@ 2 Graph state preliminaries
Figure 1: a) The square graph $G$. The associated state |$\ket{G}$, as any graph state, possesses an associated structure of the stabilizer $S_{G}$ composed of specific strings of Pauli operators. Expectation values of any Pauli string $O$ is nonzero if and only if ±$\pm O\in S_{G}$. $X_{1}X_{3}$ is a stabilizer operator of |$\ket{G}$ and therefore it’s expectation value is 1.
b) Graphs corresponding to the state defined in a) being altered by noise – here, the noise may remove certain edges (dashed lines). Noise processes change the stabilize group (e.g. by graph modification shown here), which allows for verifiable preparation.
In this section, we present a short introduction to the class of graph states and the stabilizer formalism. Readers familiar with the topic may skip directly to the next section.
A graph $G=(V,E)$ is defined by a set $V$ of vertices and a set of edges $E\subset V^{2}$: each $\{i,j\}=:e\in E$ denotes a connection between two vertices $i,j\in V$. Graph states are multi-qubit quantum states, where the vertices and edges of the graph $G=(V,E)$ represent qubits and entangling gates, respectively. The state |$\ket{G}$, corresponding to a graph $G$ is defined as
@@
where $\operatorname{CZ}$ denotes a standard controlled-$Z$ unitary operation
@@
Manifestly, all of controlled-$Z$ gates are diagonal in the computation basis, and hence commute – the product in the above equation does not depend on the order of operations.
Consider a four-vertex graph shown in Figure 1a: it has a vertex set $V=\{1,2,3,4\}$ and an edge set $E=\{\{1,2\},\{2,3\},\{3,4\},\{4,1\}\}$. The corresponding graph state is given by
|$\ket{G}=\operatorname{CZ}_{1,2}\operatorname{CZ}_{2,3}\operatorname{CZ}_{3,4}\operatorname{CZ}_{4,1}\ket{+}^{\otimes 4}.$ (3)
The graph states can also be described using the so-called stabilizer operators, providing a description of the state in question in terms of its correlation structure. Consider an initial state |$\ket{+}^{\otimes|V|}$: since it is a pure product state, there are no correlations between different sites. However, it is an eigenstate for each $X_{i}$ for $i\in V$. Let us denote the product of controlled-$Z$ operators in Equation 1 by $U$; then
@@
where $X_{i}$, $Y_{i}$, and $Z_{i}$ denote the Pauli matrices acting on the $i$-th qubit and $\operatorname{\mathcal{N}}(i)$ is the neighbourhood of $i$. As shown above, any graph state |$\ket{G}$ defined by Equation 1 is an eigenstate (to the eigenvalue of $+1$) of all $g_{i}$:
$g_{i}\ket{G}=\ket{G}$ (7)
While the $(+1)$-eigenspace of each of the operators $g_{i}$ is highly degenerated, taken together they fully define the state |$\ket{G}$: it is the $(+1)$-eigenstate of all $g_{i}$. Operators defined by Equation 7 commute; hence, they generate an Abelian subgroup of the group of Pauli operators, called the stabilizer $S$ of |$\ket{G}$:
@@
The elements of the group $S$, due to the compact form of the generators (Equation 6), can be described with the help of low-dimensional matrices of integers modulo $2$ [21]. This representation is helpful in direct calculations involving the stabilizer.
@@
The elements of $S$ form a complete description of all possible correlations of local measurements of graph states: if a Pauli string [^1] Here, $o_{i}$ denotes one of the Pauli operators $\{X_{i},Y_{i},Z_{i}\}$ acting on the $i$-th qubit. $O=o_{i_{1}}o_{i_{2}}\ldots$ (or its negative, $-O$) appears in $S$, its expectation value on |$\ket{G}$ is equal to $+1$ (respectively, $-1$), and if ±$\pm O\notin S$, $\langle O\rangle_{G}=0$.
For example, for the graph state shown in Figure 1a, the string $O=X_{1}X_{3}=g_{1}g_{3}$ is a stabilizer operator of |$\ket{G}$. Using Equation 7 it can be seen that $\langle O\rangle_{G}=1$. The string $O^{\prime}=X_{1}Z_{3}\notin S_{G}$, hence $\langle O^{\prime}\rangle_{G}=0$. In this paper we are mostly using Pauli strings which contain only Pauli $X$ operators [21]. However, the results can be generalized to other Pauli strings.
The weighted graph states [22] are a related class of quantum states, for which controlled-phase operators are used instead of controlled-$Z$, with the phases possibly different for each edge of the graph $G$. If the phase associated with the edge $\{i,j\}$ is $\varphi_{i,j}$, the resulting state is defined by a modified version of Equation 1,
@@
where $\operatorname{C\Phi}(\varphi)=\operatorname{diag}(1,1,1,\exp(i\varphi))$ on the $\{i,j\}$ subsystem.S uch states arise naturally in systems with Ising-like interactions, and offer an useful representation of noisy graph state preparation using such a scheme. Unfortunately, the analogue of procedure described by Equation 5 does not produce a Pauli string, and hence a simple stabilizer description of correlations generally does not exist.
For a standard graph state |$\ket{G}$ all phases $\varphi_{i,j}$ above are equal to $\pi$. This class of states naturally arises in systems with an Ising-type interaction pattern, and some physical realizations of graph states prepare the unweighted graph states in this exact way.
@@
|$\ket{GHZ}=\frac{1}{\sqrt{2}}(\ket{0_{1}0_{2}\dots 0_{N}}+\ket{1_{1}1_{2}\dots 1_{N}})$ (10)
is local unitary equivalent to the graph state of the fully connected $N$ vertex graph (or, equivalently, by the $N$ vertex star graph). In measurement-based quantum computing, cluster states, corresponding to grid graphs, are relevant [3], and intermediate-size cluster states have recently been realized on various platforms [24, 25]. A detailed discussion of graph states, including weighted graph states, can be found in [5, 26].
@@ 3 Correlations and measurement
It is possible to extract an entangled pair of qubits $(i,j)$ from a graph state |$\ket{G}$ by means of local Pauli measurements, if there exists a path between $i$ and $j$ through the graph $G$ [20, 3]. In general, there are multiple ways to perform this task – and if noise is present, the entanglement quality of the resulting two-qubit state depends on the choice.
Consider a graph state |$\ket{G}$ that undergoes a local sequential measurement process involving Pauli operators. The state of the unmeasured qubits is completely characterized by the measurement pattern along with the outcomes; the remaining correlations stem from the stabilizer operators of |$\ket{G}$ consistent with the measurement pattern.
To see the structure of these correlations, let us choose a measurement pattern on a subset of qubits $I$: for each qubit $i\in I$, a local Pauli $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$ is chosen. Then, each qubit $i$ is measured sequentially in the eigenbasis of the associated operator $o_{i}$. As mentioned previously, the expectation value of a Pauli string $O=\prod_{i\in I}o_{i}$ can be determined: if $O\in S_{G}$, it implies $\expectationvalue{O}_{G}=1$. The case of $-O\in S_{G}$ differs only by the sign, and if neither ±$\pm O\in S_{G}$, the expectation value is 0.
Let us concentrate on the positive sign case (the negative sign does not change much in reasoning) and consider an actual preparation of |$\ket{G}$ and subsequent measurement according to $\{o_{i}\}_{i\in V}$: the measurement of $o_{i}$ on the $i$-th qubit has an outcome of $s_{i}$. If $\langle O\rangle_{G}=1$, the outcomes of the sequential measurement must reflect that: the signs of the measurement results must multiply to $+1$. Any other outcome pattern would decrease the magnitude of the experimental expectation value and is therefore incompatible with the graph state correlations.
This reasoning applies as well if the sequential measurement is stopped at any point and resumed afterward: if $O=\prod_{i\in I}o_{i}$ is a stabilizer, and only the qubits $I^{\prime}\subset I$ were measured, the sign structure of the further measurement of $I\setminus I^{\prime}$ can be predicted. Thus, by the action of partial measurement, correlation in the remaining qubits is induced.
The following Lemma captures this line of thought: the end correlations are defined by stabilizer operators consistent with the measurement scheme. Note that the measurement scheme does not itself have to be a stabilizer operator: only part of it has to extend to one, and this is always possible (see Lemma 3 in Appendix A).
@@
Let |$\ket{G}$ be a graph state determined by the graph $G=(V,E)$; we denote the stabilizer of |$\ket{G}$ by $S$. If a subset $I$ of qubits is measured in such a way that a qubit $i\in I$ is measured in the eigenbasis of $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$, we encode this measurement scheme as the Pauli string $O=\prod_{i\in I}o_{i}$. The qubits of $I$ are measured independently and sequentially, with the measurement outcome of $o_{i}$ denoted by $s_{i}=\pm 1$.
Let us now write a stabilizer operator $Q\in S$ of |$\ket{G}$ as a Pauli string: $Q=\pm\prod_{j\in J}q_{j}$. The post-measurement state |$\ket{\psi}$ is determined by those elements of the stabilizer, which are consistent with the measurement pattern $O$ such that $Q|_{I\cap J}=O|_{I\cap J}$. The stabilizer operators of the post-measurement state |$\ket{\psi}$ can be computed from such operators $Q$ as
$Q^{\prime}:=\pm\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}q_{j}.$ (11)
where the global sign is equal to the global sign of $Q$.
@@
Consider the stabilizer operator $Q$ defined as stated in the Lemma. Then assume that the measurement outcomes of each of the qubits $i\in I\cap J$ are $s_{i}$. Then, the projective measurement operators of these qubits can be written as $\frac{1}{2}\left(\mathbbm{1}_{i}+s_{i}q_{i}\right)$. [^2] For instance, the projection operator associated with the $X+$ outcome is |$\ket{X+}\bra{X+}=\frac{1}{2}\left(\mathbbm{1}+X\right)$. Since the projection operators commute with $q_{j}$ themselves, the following chain of equalities holds:
|$\ket{\psi}$ $=C\prod_{i\in I\cap J}\frac{\mathbbm{1}_{i}+s_{i}q_{i}}{2}\underbrace{\ket{G}}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q\ket{G}}$ (12)
@@
=(±$=\underbrace{\left(\pm\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}q_{j}\right)}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q^{\prime}}\ket{\psi}.$
Here, $C$ is a normalization constant and the following equation (true for any $q$ such that $q^{2}=\mathbbm{1}$, in particular $q\in\{X,Y,Z\}$) was used:
$\frac{\mathbbm{1}+sq}{2}q=s\frac{sq+\mathbbm{1}}{2}\quad\text{ for }s=\pm 1.$ (13)
Figure 2: Partial local measurement of graph states induces correlations in the remaining qubits. Pictures is the path graph $5$ vertices: if the associated graph state is measured according to the shown measurement pattern, Bell-like correlations are induced in the two terminal qubits.
@@
As an example, consider the line graph of 5 qubits, as shown in Figure 2 and the Pauli string $O=X_{2}X_{3}X_{4}$ which has support on vertices in $I=\{2,3,4\}$. The stabilizer set of the 5 qubit line graph contains $2^{5}$ stabilizers. Three of them are $Q^{(a)}=X_{1}X_{3}X_{5}$, $Q^{(b)}=Z_{1}X_{2}X_{4}Z_{5}$, and $Q^{(c)}=-Y_{1}X_{2}X_{3}X_{4}Y_{5}$ which have support on qubits in $J^{(a)}=\{1,3,5\}$, $J^{(b)}=\{1,2,4,5\}$, and $J^{(c)}=\{1,2,3,4,5\}$, respectively. It holds that $Q^{(a)}|_{3}=O|_{3}$, $Q^{(b)}|_{2,4}=O|_{2,4}$, and $Q^{(c)}|_{2,3,4}=O|_{2,3,4}$. Following Lemma 1, the post-measurement state after measuring qubits 2, 3, 4 in the $X$-basis is stabilized by $s_{3}X_{1}X_{5}$, $s_{2}s_{4}Z_{1}Z_{5}$, and $-s_{2}s_{3}s_{4}Y_{1}Y_{5}$. This can be viewed as a graph state (Bell pair) in a nonstandard basis – the change is an effect of the measurement performed on the qubits $\{2,3,4\}$.

### v2

@@ 1 Introduction
The discovery of cluster states – states of qubits with grid-like entanglement structure – provided a new perspective for quantum computation, better applicable for some systems [1, 2]. The original design consisted of qubits arranged in a two-dimensional square lattice structure, where the neighboring qubits would be entangled by means of a controlled-$Z$ gates. Unlike the circuit-based understanding of quantum computation [3, 4], the input state is constant, and the only available class of operations is the sequential measurement of individual qubits. It is the act of measurement that feeds the information into the system and performs all of the needed transformations [5]. Generalizations into different connectivity structures – the graph and hypergraph states – soon followed, sharing the same core properties with the original cluster states [6, 7, 8, 9] and providing a framework for further research into the entanglement properties of multiqubit states [10, 11]. Another task for which graph states are useful is the distribution of entanglement in quantum networks [12, 13, 14, 15, 16, 17].
@@
In the context of quantum computation, the most basic operation is no processing at all: a simple transfer or teleportation [18, 19] of one qubit state into another place. It is a basic building block of measurement-based quantum computation, on top of which more involved operations are built. In the basic scenario, teleportation can be thought of as a creation of a Bell pair, which is subsequently used to transmit the quantum state. The Bell pair creation step can be performed by sequential measurement of Pauli-$X$ operators on a path between the initial and end qubits and Pauli-$Z$ on the qubits neighboring the path (see Figure 1). Intuitively, the $Z$-basis measurement excises the entanglement contained in the path from the rest of the cluster state and the $X$-basis measurement along the path leaves the terminal (first and last) qubits in Bell-type correlations, up to local unitary rotations. The final measurement of one of them fuses the two together, and the surviving qubit shares correlations with a distant part of the graph ([5], see also Appendix A).
The crucial point is that the Bell pair creation step is independent on the state to be transferred. The right correlational structure is prepared and subsequently used, and the preparation step can be studied alone. This abstraction step is helpful in various ways: the theoretical effects of the exact input state for teleportation are not relevant, this formulation allows us to utilize the works dedicated to quantum state extraction [20, 21], and similar language regarding the correlation of the relevant qubits is frequently used [22, 23], also in the research dedicated to teleportation [24].
Figure 1: Any two vertices of a graph state associated to a connected graph can be turned into a Bell pair by sequential measurements. a) A grid graph describing the underlying structure of a cluster state. The goal is to generate a Bell pair between the qubits associated to the big red vertices. The chosen path is highlighted in red. b) We measure the qubits associated with the vertices on the path between the red vertices in $X$ basis and the qubits adjacent to the chosen path in $Z$ basis. c) The post-measurement state has Bell correlations between the qubits associated to the red vertices. This procedure was called X-protocol in [12].
@@
In general, entanglement is required for quantum computation, and preparation of a Bell pair or GHZ state is an operational primitive of many quantum algorithms. Extracting those states from larger graph states was discussed in [25, 26, 24, 13, 12]. The problem of extracting graph states from larger graph states is known to be NP-complete [27, 28].
@@
In this article we analyze experimentally motivated noise models and determine what can be done to the initial noisy graph state structure and the measurement pattern in order to mitigate the noise effects. This general problem of noise reduction in graph states has been analyzed from various viewpoints. For instance, purification and distillation protocols are discussed in [29, 30, 31]. Quantum error correcting codes are more general, but complex. Unlike them, our approach utilizes only local single qubit measurements, basic preparation and verification, and is in principle adaptable to various underlying graph structures. However, it still uses the properties of the stabilizer associated with the graph state. In the ideal case, some outcomes are perfectly (anti)correlated, and any deviation indicates the presence of noise. As mentioned above, teleportation is equivalent to a Bell pair preparation; in order to simplify the calculations, we analyze the entanglement quality of the prepared state after the internal qubits have been measured.
@@
The article is structured as follows. In Section 2, we briefly introduce graph states and the stabilizer formalism. In Section 3, we explain the measurement scheme which we use to generate Bell pairs between certain nodes or for teleportation. We introduce noise models in Section 4 and analyze how to deal with their effects in Section 5. In Section 6, we show how to extend the presented methods to generate GHZ states. We conclude our work in Section 7. A portion of this article contains numerical results, the code [32] used in data generation and analysis is publicly available.
@@ 2 Graph state preliminaries
Figure 2: a) The square graph defines the associated graph state $\ket{\square}$ which, as any graph state, possesses an associated structure of the stabilizer $S_{\square}$ composed of specific strings of Pauli operators. Expectation values of any Pauli string $O$ is nonzero if and only if one of ±$\pm O$ belongs to $S_{\square}$. $X_{1}X_{3}$ is a stabilizer operator of $\ket{\square}$ and therefore it’s expectation value is 1.
b) Graph states corresponding to the state defined in a) being altered by noise – here, the noise may remove certain edges (dashed lines), as in the exemplary $\ket{\sqcap}$ defined by the graph shown in the top left corner. Noise processes change the stabilizer group (e.g. by graph modification shown here), which allows for verifiable preparation.
In this section, we present a short introduction to the class of graph states and the stabilizer formalism.
A graph $G=(V,E)$ is defined by a set $V$ of vertices and a set of edges $E\subset V^{2}$: each $\{i,j\}=:e\in E$ denotes a connection between two vertices $i,j\in V$ , where $i\neq j$. Graph states are multi-qubit quantum states, where the vertices and edges of the graph $G=(V,E)$ represent qubits and entangling gates, respectively. The state |$\ket{G}$, corresponding to a graph $G$, is defined as
@@
where $\operatorname{CZ}$ denotes a standard controlled-$Z$ unitary operation on qubits $i$ and $j$
@@
The controlled-$Z$ gates are diagonal in the computation basis, and hence commute. Therefore, the product in the above equation does not depend on the order of operations.
Consider a four-vertex square graph shown in Figure 2a: it has a vertex set $V^{\square}=\{1,2,3,4\}$ and an edge set $E^{\square}=\{\{1,2\},\{2,3\},\{3,4\},\{4,1\}\}$. The corresponding graph state is given by
|$\ket{\square}=\operatorname{CZ}_{1,2}\operatorname{CZ}_{2,3}\operatorname{CZ}_{3,4}\operatorname{CZ}_{4,1}\ket{+}^{\otimes 4}.$ (3)
The graph states can also be described using the so-called stabilizer operators, providing a description of the state in question in terms of its correlation structure. Consider an initial state |$\ket{+}^{\otimes|V|}$: since it is a pure product state, there are no correlations between different sites. However, it is a $(+1)$-eigenstate for each $X_{i}$ for $i\in V$. Let us denote the product of controlled-$Z$ operators in Equation 1 by $U$; then
@@
where $X_{i}$, $Y_{i}$, and $Z_{i}$ denote the Pauli matrices acting on the $i$-th qubit and $\operatorname{\mathcal{N}}(i)$ is the neighborhood of the vertex $i$, given by
$\operatorname{\mathcal{N}}(i)=\{j:\{i,j\}\in E\}.$ (7)
As shown above, any graph state |$\ket{G}$ defined by Equation 1 is an eigenstate with eigenvalue $+1$ of all $g_{i}$:
$g_{i}\ket{G}=\ket{G}.$ (8)
While the $(+1)$-eigenspace of each of the operators $g_{i}$ is highly degenerated, taken together they fully define the state |$\ket{G}$. The graph state |$\ket{G}$ is the unique $(+1)$-eigenstate of all $g_{i}$. Operators defined by Equation 8 commute; hence, they generate an Abelian subgroup of the group of Pauli operators, called the stabilizer $S$ of |$\ket{G}$:
@@
The elements of the group $S$, due to the compact form of the generators (Equation 6), can be described with the help of low-dimensional matrices of integers modulo $2$ [33]. This representation is helpful in direct calculations involving the stabilizer group.
@@
The elements of the stabilizer $S$ form a complete description of all possible correlations of local measurements of graph states. Let $O=o_{i_{1}}o_{i_{2}}\ldots$ be a Pauli string, where $o_{i}$ denotes one of the Pauli operators $\{X_{i},Y_{i},Z_{i}\}$ acting on the $i$-th qubit. If ±$\pm O$ appears in $S$, its expectation value on |$\ket{G}$ is equal to ±$\pm 1$. Otherwise, if ±$\pm O\notin S$, we have $\langle O\rangle_{G}=0$.
For example, for the graph state shown in Figure 2a, the string $O=X_{1}X_{3}=g_{1}g_{3}$ is a stabilizer operator of $\ket{\square}$. Using Equation 8 it can be seen that $\langle O\rangle_{\square}=1$. The string $O^{\prime}=X_{1}Z_{3}\notin S_{\square}$, hence $\langle O^{\prime}\rangle_{\square}=0$. In this paper we are mostly using Pauli strings which contain only Pauli $X$ operators [33]. However, the results can be generalized to other Pauli strings.
Weighted graph states [34] are a related class of quantum states, for which controlled-phase operators are used instead of controlled-$Z$. The phases $\varphi_{i,j}$ are possibly different for each edge $\{i,j\}\in E$ of the graph $G_{\varphi}=(V,E,(\varphi_{i,j}))$. A modified version of Equation 1 which describes the weighted graph state is
@@
where $\operatorname{C\Phi}(\varphi)=\operatorname{diag}(1,1,1,\exp(i\varphi))$ on the $\{i,j\}$ subsystem, similarly to Equation 2. Such states arise naturally in systems with Ising-like interactions, and offer an useful representation of noisy graph state preparation using such a scheme. Unfortunately, the analogue of procedure described by Equation 5 does not produce a Pauli string, and hence a simple stabilizer description of correlations generally does not exist. For a standard graph state |$\ket{G}$ all phases $\varphi_{i,j}$ above are equal to $\pi$.
@@
|$\ket{GHZ}=\frac{1}{\sqrt{2}}(\ket{0_{1}0_{2}\dots 0_{N}}+\ket{1_{1}1_{2}\dots 1_{N}}),$ (11)
is local unitary equivalent to the graph state of the fully connected $N$ vertex graph or, equivalently, by the $N$ vertex star graph. In measurement-based quantum computing the cluster states, corresponding to grid graphs, are relevant [5] and are an example of a broader category of universal resource states [36]. Intermediate-size cluster states have recently been realized on various platforms [37, 38]. A detailed discussion of graph states, including weighted graph states, can be found in [7, 39].
@@ 3 Correlations and measurement
In this section we show how local measurement affects correlations present in graph states. In particular, it is possible to extract an entangled pair of qubits $(i,j)$ from a graph state |$\ket{G}$ by means of local Pauli measurements, if there exists a path between $i$ and $j$ through the graph $G$ [12, 5]. As we later show in Section 5, the basic procedure outlined in Figure 3 is susceptible to noise effects, and the general mathematical language introduced here is used in the analysis and proposal of alternative extraction procedures.
The post-measurement state after applying a sequence of Pauli measurements can be determined by applying graphical rules for measurements [7]. In this section, we present an equivalent method using stabilizer operators to find the post-measurement states. This method is better suited for calculations. In general, there are multiple choices of measurement sequences. If noise is present, the entanglement quality of the resulting two-qubit state depends on the choice of the sequence.
Consider a graph state |$\ket{G}$ that undergoes a local sequential measurement process involving Pauli operators.
The state of the unmeasured qubits is completely characterized by the measurement pattern, that is, the choice of sequential Pauli measurements, along with the outcomes. The remaining correlations stem from the stabilizer operators of |$\ket{G}$ consistent with the measurement pattern.
To see the structure of these correlations, let us choose a measurement pattern on a subset of qubits $I$: for each qubit $i\in I$, a local Pauli $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$ is chosen. Then, each qubit $i$ is measured sequentially in the eigenbasis of the associated operator $o_{i}$. As mentioned previously, the expectation value of a Pauli string $O=\prod_{i\in I}o_{i}$ can be determined: if $O\in S_{G}$, it implies $\expectationvalue{O}_{G}=1$. The case of $-O\in S_{G}$ differs only by the sign, and if neither ±$\pm O$ belongs to $S_{G}$, the expectation value is 0.
Without loss of generality, let us concentrate on the positive sign case and consider a preparation of |$\ket{G}$ and subsequent measurement according to $\{o_{i}\}_{i\in V}$: the measurement of $o_{i}$ on the $i$-th qubit has an outcome of $s_{i}{=\pm 1}$. If $\langle O\rangle_{G}=1$, the outcomes of the sequential measurement must reflect that: the signs of the measurement results must multiply to $+1$. Any other outcome pattern would decrease the magnitude of the expectation value (defined empirically as average over random sequence of results) and is therefore incompatible with the graph state correlations.
This reasoning applies as well if the sequential measurement is stopped at any point and resumed afterward. If $O=\prod_{i\in I}o_{i}$ is a stabilizer operator, and only the qubits $I^{\prime}\subset I$ were measured, the sign structure of the further measurement of $I\setminus I^{\prime}$ can be predicted. Thus, by the action of partial measurement, correlation in the remaining qubits is induced.
The following Lemma captures this line of thought: the end correlations are defined by stabilizer operators consistent with the measurement scheme. Note that the measurement scheme does not itself have to be a stabilizer operator: only part of it has to extend to one, and such an extension always exists (see Lemma 4 in Appendix B).
@@
Let |$\ket{G}$ be a graph state determined by the graph $G=(V,E)$. We denote the stabilizer of |$\ket{G}$ by $S$. If a subset $I$ of qubits is measured in such a way that a qubit $i\in I$ is measured in the eigenbasis of $o_{i}\in\{X_{i},Y_{i},Z_{i}\}$, we encode this measurement scheme as the Pauli string $O=\prod_{i\in I}o_{i}$. The qubits of $I$ are measured independently and sequentially, with the measurement outcome of $o_{i}$ denoted by $s_{i}=\pm 1$.
Let us now write a stabilizer operator $Q\in S$ of |$\ket{G}$ as a Pauli string with support on indices $J$: $Q=\pm\prod_{j\in J}q_{j}$, where $q_{j}$ is a Pauli operator in $\{X_{j},Y_{j},Z_{j}\}$, acting on qubit $j\in J$. The post-measurement state |$\ket{\psi}$ is determined by those elements of the stabilizer, which are consistent with the measurement pattern $O$ such that $Q|_{I\cap J}=O|_{I\cap J}$. The stabilizer operators of the post-measurement state |$\ket{\psi}$ can be computed from such operators $Q$ as
$Q^{\prime}:=\pm\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}q_{j},$ (12)
where the global sign $\pm$ in front of the product is equal to the the one appearing in the definition of $Q$.
@@
Consider the stabilizer operator $Q$ defined as stated in the Lemma. Then assume that the measurement outcomes of each of the qubits $i\in I\cap J$ is $s_{i}$. The projective measurement operators of these qubits can be written as $\frac{1}{2}\left(\mathbbm{1}_{i}+s_{i}q_{i}\right)$. For instance, the projection operator associated with the $X+$ outcome is |$\outerproduct{X+}{X+}=\frac{1}{2}\left(\mathbbm{1}+X\right)$. Since the projection operators commute with $q_{j}$ themselves, the following chain of equalities holds:
|$\ket{\psi}$ $=C\prod_{i\in I\cap J}\frac{\mathbbm{1}_{i}+s_{i}q_{i}}{2}\!\!\!\underbrace{\ket{G}}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q\ket{G}}$
@@
=(±$=\underbrace{\left(\pm\!\!\!\prod_{i\in I\cap J}s_{i}\prod_{j\in J\setminus I}\!\!\!q_{j}\right)}_{\leavevmode\nobreak\ \leavevmode\nobreak\ =Q^{\prime}}\ket{\psi}.$ (13)
Here, $C$ is a normalization constant and the following property
$\frac{\mathbbm{1}+sq}{2}q=s\frac{sq+\mathbbm{1}}{2}\quad\text{ for }s=\pm 1$ (14)
was used, valid for any $q$ such that $q^{2}=\mathbbm{1}$ and in particular for $q\in\{X,Y,Z\}$. ∎
Figure 3: Partial local measurements on a graph states induces correlations in the remaining qubits. a) If the 4 vertex path graph state is measured according to the shown measurement pattern, that is, measuring the inner vertices in the $X$-basis, we get a connected two-vertex graph. b) The graph here is not connected: the edge $\{2,3\}$ is absent, possibly due to scenario outlined in Section 4.1. If the corresponding graph state is measured according to the same measurement pattern, we get a disconnected two-vertex graph. Measurements can never induce entanglement between parties that were separable before.
@@
As an example, consider the path graph of 4 qubits, as shown in Figure 3a and the Pauli string $O=X_{2}X_{3}$ which has support on the vertices in $I=\{2,3\}$. The stabilizer of the four-qubit path graph contains $2^{4}$ stabilizer elements. Three of them are $Q^{(a)}=X_{1}X_{3}Z_{4}$, $Q^{(b)}=Z_{1}X_{2}X_{4}$, and $Q^{(c)}=-Y_{1}X_{2}X_{3}Y_{4}$ which have support on qubits in $J^{(a)}=\{1,3,4\}$, $J^{(b)}=\{1,2,4\}$, and $J^{(c)}=\{1,2,3,4\}$, respectively. We have $Q^{(a)}|_{3}=O|_{3}$, $Q^{(b)}|_{2}=O|_{2}$, and $Q^{(c)}|_{2,3}=O|_{2,3}$. [^1] The notation $O|_{S}$ denotes restriction of a product of operators to a qubit subset $S$. Following Lemma 1, the post-measurement state after measuring qubits 2 and 3 in the $X$-basis is stabilized by $s_{3}X_{1}Z_{4}$, $s_{2}Z_{1}X_{4}$, and $-s_{2}s_{3}Y_{1}Y_{4}$. For $s_{2},s_{3}=+1$, these are stabilizer operators of the two-vertex graph state with one edge between them, as shown in the lower part of Figure 3a. For other measurement outcomes $s_{2},s_{3}$, we get stabilizer operators which are equivalent up to local unitary operations. Recall that this graph state is local unitary equivalent to the Bell states.
Another example of a four-vertex disconnected graph is shown in Figure 3b. This graph state is also stabilized by the $2^{4}$ operators. The three of them that coincide with the measurement pattern $O=X_{2}X_{3}$ are $Q^{(a)}=Z_{1}X_{2}$, $Q^{(b)}=X_{3}Z_{4}$, and $Q^{(c)}=Z_{1}X_{2}X_{3}Z_{4}$ that have support in $J^{(a)}=\{1,2\}$, $J^{(b)}=\{3,4\}$, and $J^{(c)}=\{1,2,3,4\}$, respectively. We have $Q^{(a)}|_{2}=O|_{2}$, $Q^{(b)}|_{3}=O|_{3}$, and $Q^{(c)}|_{2,3}=O|_{2,3}$. Following Lemma 1, the post-measurement state after measuring qubits 2 and 3 in the $X$-basis is stabilized by $s_{2}Z_{1}\operatorname{\mathds{1}}_{4}$, $s_{3}\operatorname{\mathds{1}}_{1}Z_{4}$, and $s_{2}s_{3}Z_{1}Z_{4}$. No correlations are present, since both remaining qubits are stabilized by independent local operators, and the resulting state is a product one.
Note that the stabilizer operators of post-measurement states are not always graph state stabilizer operators, but are always locally unitarily equivalent to them [7]. The example in Figure 3a is chosen so that the post-measurement state does not require this local unitary transformations. Choosing a path graph of odd length results in different Bell state stabilizers. In the example shown in Figure 3b, we get compositions of Pauli $Z$-operators as post-measurement stabilizers. Applying the local unitary $H_{1}H_{4}$ (where $H$ is the Hadamard gate) to them leads to the actual graph state stabilizer operators.